# A Survey of Reinforcement Learning For Economics

### 23 Mar 2026 v4

@@ 6.5 Implementation Dominates Algorithmic Innovation
Engstrom et al. (2020) find that PPO with the clipping mechanism disabled performs indistinguishably from the full algorithm; TRPO (Schulman et al., 2015) with the same code-level additions matches PPO.120120120The non-clipping components that suffice are: observation normalization, reward normalization, value function clipping, global gradient norm clipping, orthogonal weight initialization, and the Adam optimizer. Andrychowicz et al. (2021) identify observation normalization, orthogonal initialization, and learning rate annealing as the three choices accounting for most variance across 250,000 agents and 250 hyperparameter configurations. Huang et al. (2022) catalog 37 implementation details required to reproduce PPO on Atari;121121121These include reward clipping to [1,1][-1,1], frame stacking to 4 frames, a specific episode termination convention, and a numerically stable normalization of the advantage estimate. omitting any produces materially different results.
The most consequential implementation detail is the distinction between termination and truncation (Pardo et al., 2018). In reinforcement learning, an episode can end for two distinct reasons: termination, where the environment reaches a natural absorbing state (the pole falls in CartPole, the robot falls in locomotion tasks), and truncation, where the episode is cut short by an external time limit. At a termination, the value of the successor state is zero: $V(s_{\text{term}})=0$. At a truncation, the episode is merely paused and the successor state has non-zero value: $V(s_{\text{trunc}})\neq 0$. Treating truncated transitions as terminated substitutes zero for a non-zero bootstrap value at every time limit boundary, corrupting every Bellman update in the vicinity of episode boundaries. Pardo et al. (2018) show that this conflation degrades performance by 20–40% on standard MuJoCo benchmarks.122122122The Gymnasium API (Towers et al., 2024) enforces the distinction by returning separate terminated and truncated flags, but most pre-2022 codebases conflate them in the done flag.
@@ 6.6 Replay Buffer Pathologies and Reward Scaling
Experience replay (Lin, 1992) decouples the data collection and learning processes, allowing a single transition to be used for multiple gradient updates. The replay ratio (the number of gradient updates per environment step) governs the trade-off between sample efficiency and data staleness. Zhang and Sutton (2017) show that increasing the replay ratio beyond a modest threshold degrades performance.123123123At high replay ratios, the distribution shift between the current policy and the behavior policy that generated the stored data grows large enough to violate the off-policy assumptions of Q-learning. The relationship is non-monotone and environment-dependent, making replay buffer size a sensitive hyperparameter with no universal default.
Schaul et al. (2016) propose prioritized experience replay (PER).124124124PER samples transitions with probability proportional to the magnitude of their TD error, on the argument that high-error transitions are the most informative. Fedus et al. (2020) revisit PER on large-scale Atari experiments and find that uniform sampling from a large enough buffer matches or outperforms PER, while being simpler to implement and tune.
Reward scaling introduces a separate class of failure modes. Standard DQN (Mnih et al., 2015) clips rewards to [1,+1][-1,+1] across all environments to stabilize training. van Hasselt et al. (2016a) observe that reward clipping changes the objective: clipped rewards make all positive events equivalent regardless of magnitude, so the agent learns to maximize the frequency of positive events rather than their cumulative value. This substitution can produce policies that are locally rational under the clipped reward but qualitatively suboptimal under the true reward. van Hasselt et al. (2016a) propose PopArt as a remedy.125125125PopArt normalizes targets to have unit variance while adjusting the output layer so that the policy remains invariant to the normalization. PopArt allows consistent learning across reward scales spanning several orders of magnitude without reward clipping.
Skalse et al. (2022) formalize reward hacking and show that any non-constant proxy reward can in principle be exploited by a sufficiently capable optimizer.126126126Skalse et al. (2022) define a proxy as unhackable if increasing expected proxy return cannot decrease expected true return. Their main result states that for the set of all stochastic policies, two reward functions are unhackable only if one of them is constant. For deterministic policies and finite policy sets, non-trivial unhackable pairs exist, but the conditions are stringent.
@@ 6.7 Simulation Study: Bellman Error and Value Error in Offline Policy Evaluation
Two algorithms are trained on $\mathcal{D}$.12712712750,000 gradient steps per seed, 3 seeds; two-layer MLP with 64 hidden units, Adam at $5\times 10^{-4}$; target network updated every 500 steps. BRM minimizes $\mathbb{E}_{\mathcal{D}}[(Q(s,a)-(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})))^{2}]$ where both $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ use the current network; gradients flow through both sides simultaneously. The key consequence is that the network can zero the residual at an observed pair $(s,a)$ by co-moving $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ together, rather than moving either toward $Q^{*}$. This is the opposite of supervised learning, where labels are fixed external targets that do not move with the model weights. FQE prevents this by using a frozen target network for $Q(s^{\prime},a^{\prime})$; the network must reduce $Q(s,a)$ toward a target that does not respond to its own gradient steps. Every 500 steps we record the Bellman error on $\mathcal{D}$ (current network on both sides) and the value error $\frac{1}{|\mathcal{F}|}\sum_{(s,a)\in\mathcal{F}}|Q_{\theta}(s,a)-Q^{*}(s,a)|$ over all 4,795 feasible pairs. As a supervised baseline, OLS regression of $\log c$ on $\log k$ and $\log z$ is estimated on expanding windows.128128128Noise $\sigma=0.30$; windows from $n=10$ to 2,0002{,}000; held-out test set of 500 transitions.
@@ 7 Reinforcement Learning for Optimal Control
A handful of organizations have deployed reinforcement learning beyond simulation, achieving measurable gains on specific large-scale problems. Each deployment required substantial domain engineering and scientific tuning; these remain exceptional cases rather than standard practice. I review the most prominent field deployments, including ride-hailing dispatch at DiDi, data center cooling at Google, hotel revenue management, and financial order execution, before concluding with a simulation study on the bus engine replacement problem. In each case, the RL agent’s parameters were updated during a training phase conducted in simulation or on historical data, and the resulting policy was deployed with fixed weights (Section 2).
@@ 7.1 Ride-Hailing Dispatch
Qin et al. (2021) formalized DiDi’s order dispatching as a semi-Markov decision process129129129A semi-Markov decision process generalizes the standard MDP by allowing variable time between decisions. The discount factor $\gamma^{\tau_{k}}$ depends on the actual time elapsed $\tau_{k}$ rather than applying a fixed per-step discount. where each driver is an independent agent. The state of a driver consists of location (discretized into hexagonal zones) and time (bucketed into intervals). The action is the order assigned to the driver, with the option to remain idle. The reward is the trip fare. State transitions are determined by trip destinations, as completing an order transports the driver from origin to destination, changing their spatial state. The stochasticity arises from future demand, which determines the available actions at each state. The per-driver value function $V^{\pi}(s)$ represents the expected cumulative fare a driver can earn from state $s$ under dispatching policy $\pi$:
@@
where $\hat{r}_{i}$ is the predicted fare for order $i$, $\hat{\tau}_{i}$ is the estimated trip duration, and $s_{j}^{\prime}$ is the destination state. The value function is learned via temporal-difference methods from historical trip data. DiDi’s Cerebellar Value Network (CVNet; Tang et al., 2019) uses hierarchical coarse-coding130130130Coarse-coding represents a value function as a weighted sum over overlapping rectangular or hexagonal tiles that partition the state space (Sutton and Barto, 2018); each state activates the tiles that contain it, and a hierarchical variant stacks tiles at multiple resolutions to allow both coarse and fine generalization simultaneously. with a multi-resolution hexagonal grid, enabling transfer learning131131131Transfer learning initializes a model for a new task (a new city) with parameters trained on related tasks (existing cities), on the assumption that learned representations of traffic and demand patterns generalize across contexts and require less data to reach good performance in the new setting. across cities and robustness to data sparsity in low-traffic zones.
@@
Han et al. (2022) reported complementary results from Lyft, where the dispatching system optimizes driver assignment and repositioning jointly. Their value decomposition architecture132132132Value decomposition decomposes the platform’s global matching objective into individual driver value functions, each estimable via temporal-difference learning. The global optimum is recovered by optimizing the sum of individual values. assigns credit to individual driver decisions within the global objective. The production system demonstrated improvements in rider wait times and driver utilization, providing a second data point suggesting that similar approaches may transfer across platforms.
@@ 7.2 Hotel Revenue Management
Their system uses a two-step design. The RL agent observes the state $(t,s)$, where $t$ indexes the booking period within a $T=10$-period episode and $s$ is the average revenue per room sold to date, and selects an average discount level $a\in\{10\%,20\%,30\%\}$. A linear program then converts this scalar recommendation into a feasible capacity allocation across rate segments, accounting for each hotel’s channel preferences.133133133The RL component uses a modified on-policy Monte Carlo method with $\varepsilon$-greedy exploration, updating Q-values from realized returns after each completed episode. Training on real operational data rather than simulated environments means the method adapts to demand non-stationarity without requiring an explicit demand model. This decomposition addresses practitioner acceptance, since managers understand a single discount recommendation, and circumvents the need to explicitly model demand arrivals or cancellation rates.
The field experiment randomly assigned five Hanting-brand hotels in Shanghai to the treatment group from ten candidates, with 271 additional Shanghai hotels serving as a donor pool for synthetic control estimation.134134134Synthetic control constructs a weighted combination of untreated hotels whose pre-treatment trend matches each treated hotel, avoiding the selection bias of simple before-after comparisons. Table 10 reports the average treatment effects over the pilot period (March–June 2015).
@@ 7.3 E-Commerce Dynamic Pricing
where $\text{uv}_{i,t}$ is unique visitors in period $t$ and $\tau$ is a reference lag.135135135DRCR normalizes revenue by traffic to remove demand fluctuations unrelated to pricing. The differencing further removes product-specific level effects, yielding a reward signal that is more concave than raw revenue and improves convergence stability.
Two algorithms are deployed: DQN for the discrete action formulation and DDPG for continuous actions. Both are pre-trained from demonstrations using historical specialist pricing actions (DQfD, DDPGfD) to address cold-start.136136136Pre-populating replay buffers with rollouts from heuristic or scripted policies is a general technique for reducing exploration burden in deep RL. In robotic grasping, Kalashnikov et al. (2018) collected 200,000 scripted grasping attempts at 15–30% success rate, sufficient to bootstrap a vision-based policy to 96% success; see Ibarz et al. (2021) for a survey of these and related warm-start strategies.137137137Standard A/B testing is infeasible because Chinese e-commerce regulations prohibit displaying different prices to different customers for the same product simultaneously. Liu et al. (2019) instead evaluate using difference-in-differences against “simi-products” (similar products not subject to algorithmic pricing) as controls.
@@ 7.4 Financial Order Execution
Nevmyvaka et al. (2006) applied tabular Q-learning to execution on real limit order book data from NASDAQ stocks.138138138The dataset covers 500 trading days of millisecond-level limit order book snapshots for AMZN, QCOM, and NVDA. The state space contains approximately 10,000 states; the horizon is $T=60$ seconds with discount $\gamma=1$. The state at time $t$ is $s_{t}=(t,q_{t},\psi_{t},\Delta_{t})$, where $t\in\{1,\ldots,T\}$ is time remaining, $q_{t}\in\{0,1,\ldots,Q\}$ is inventory remaining, $\psi_{t}$ is the discretized bid-ask spread, and $\Delta_{t}$ is the discretized signed volume imbalance.139139139Signed volume imbalance is the difference between buy and sell order volume near the best prices, normalized by total volume; positive imbalance typically predicts short-term price increases. Actions $a_{t}\in\{0,\delta,2\delta,\ldots,q_{t}\}$ specify shares to execute in the current interval. The reward is the negative per-period slippage contribution:
@@
where $p_{0}$ is the mid-quote at order arrival and $p_{t}^{\text{exec}}$ is the average execution price. Total implementation shortfall is $IS=-\sum_{t=1}^{T}r_{t}/Q$.140140140Implementation shortfall measures the total cost of executing a trade relative to the benchmark mid-quote at arrival. It captures market impact, timing cost, and opportunity cost of unexecuted shares. Because the objective is cost minimization, the Q-learning update uses $\min$ over next-period actions:
@@ 7.5 Supply Chain Inventory Management
Multi-echelon inventory systems coordinate ordering decisions across supply chain stages, where each stage’s order becomes the next stage’s incoming shipment. A retailer orders from a warehouse, which orders from a distribution center, which orders from a factory. The sequential nature creates complex dependencies, since an order placed upstream today affects downstream availability many periods later. The state space grows exponentially in the number of echelons, with $K$ stages each having $M$ possible inventory levels yielding $M^{K}$ states. Classical inventory theory provides elegant solutions for special cases, most notably the echelon base-stock policy of Clark and Scarf (1960).
@@
Gijsbrechts et al. (2022) conducted a systematic evaluation of deep RL against classical base-stock policies across lost sales, dual sourcing, and multi-echelon configurations. The classical benchmark is the echelon base-stock policy, in which each stage $k$ maintains an echelon inventory position and orders to bring this position to a target level $S_{k}$.141141141Formally, the echelon inventory position at stage $k$ equals on-hand inventory at $k$ plus all inventory at downstream stages $1,\ldots,k-1$ plus in-transit inventory, minus backorders. The echelon base-stock policy of Clark and Scarf (1960) is optimal for serial systems with linear costs and backorders. For single-echelon systems with backorders, the optimal base-stock level is given by the newsvendor critical fractile $S^{*}=F_{D}^{-1}(b/(b+h))$. Table 13 summarizes results from their multi-echelon experiments.

### 24 Mar 2026 v5

@@ 6.5 Implementation Dominates Algorithmic Innovation
Engstrom et al. (2020) find that PPO with the clipping mechanism disabled performs indistinguishably from the full algorithm; TRPO (Schulman et al., 2015) with the same code-level additions matches PPO.122122122The non-clipping components that suffice are: observation normalization, reward normalization, value function clipping, global gradient norm clipping, orthogonal weight initialization, and the Adam optimizer. Andrychowicz et al. (2021) identify observation normalization, orthogonal initialization, and learning rate annealing as the three choices accounting for most variance across 250,000 agents and 250 hyperparameter configurations. Huang et al. (2022) catalog 37 implementation details required to reproduce PPO on Atari;123123123These include reward clipping to [1,1][-1,1], frame stacking to 4 frames, a specific episode termination convention, and a numerically stable normalization of the advantage estimate. omitting any produces materially different results.
The most consequential implementation detail is the distinction between termination and truncation (Pardo et al., 2018). In reinforcement learning, an episode can end for two distinct reasons: termination, where the environment reaches a natural absorbing state (the pole falls in CartPole, the robot falls in locomotion tasks), and truncation, where the episode is cut short by an external time limit. At a termination, the value of the successor state is zero: $V(s_{\text{term}})=0$. At a truncation, the episode is merely paused and the successor state has non-zero value: $V(s_{\text{trunc}})\neq 0$. Treating truncated transitions as terminated substitutes zero for a non-zero bootstrap value at every time limit boundary, corrupting every Bellman update in the vicinity of episode boundaries. Pardo et al. (2018) show that this conflation degrades performance by 20–40% on standard MuJoCo benchmarks.124124124The Gymnasium API (Towers et al., 2024) enforces the distinction by returning separate terminated and truncated flags, but most pre-2022 codebases conflate them in the done flag.
@@ 6.6 Replay Buffer Pathologies and Reward Scaling
Experience replay (Lin, 1992) decouples the data collection and learning processes, allowing a single transition to be used for multiple gradient updates. The replay ratio (the number of gradient updates per environment step) governs the trade-off between sample efficiency and data staleness. Zhang and Sutton (2017) show that increasing the replay ratio beyond a modest threshold degrades performance.125125125At high replay ratios, the distribution shift between the current policy and the behavior policy that generated the stored data grows large enough to violate the off-policy assumptions of Q-learning. The relationship is non-monotone and environment-dependent, making replay buffer size a sensitive hyperparameter with no universal default.
Schaul et al. (2016) propose prioritized experience replay (PER).126126126PER samples transitions with probability proportional to the magnitude of their TD error, on the argument that high-error transitions are the most informative. Fedus et al. (2020) revisit PER on large-scale Atari experiments and find that uniform sampling from a large enough buffer matches or outperforms PER, while being simpler to implement and tune.
Reward scaling introduces a separate class of failure modes. Standard DQN (Mnih et al., 2015) clips rewards to [1,+1][-1,+1] across all environments to stabilize training. van Hasselt et al. (2016a) observe that reward clipping changes the objective: clipped rewards make all positive events equivalent regardless of magnitude, so the agent learns to maximize the frequency of positive events rather than their cumulative value. This substitution can produce policies that are locally rational under the clipped reward but qualitatively suboptimal under the true reward. van Hasselt et al. (2016a) propose PopArt as a remedy.127127127PopArt normalizes targets to have unit variance while adjusting the output layer so that the policy remains invariant to the normalization. PopArt allows consistent learning across reward scales spanning several orders of magnitude without reward clipping.
Skalse et al. (2022) formalize reward hacking and show that any non-constant proxy reward can in principle be exploited by a sufficiently capable optimizer.128128128Skalse et al. (2022) define a proxy as unhackable if increasing expected proxy return cannot decrease expected true return. Their main result states that for the set of all stochastic policies, two reward functions are unhackable only if one of them is constant. For deterministic policies and finite policy sets, non-trivial unhackable pairs exist, but the conditions are stringent.
@@ 6.7 Simulation Study: Bellman Error and Value Error in Offline Policy Evaluation
Two algorithms are trained on $\mathcal{D}$.12912912950,000 gradient steps per seed, 3 seeds; two-layer MLP with 64 hidden units, Adam at $5\times 10^{-4}$; target network updated every 500 steps. BRM minimizes $\mathbb{E}_{\mathcal{D}}[(Q(s,a)-(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})))^{2}]$ where both $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ use the current network; gradients flow through both sides simultaneously. The key consequence is that the network can zero the residual at an observed pair $(s,a)$ by co-moving $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ together, rather than moving either toward $Q^{*}$. This is the opposite of supervised learning, where labels are fixed external targets that do not move with the model weights. FQE prevents this by using a frozen target network for $Q(s^{\prime},a^{\prime})$; the network must reduce $Q(s,a)$ toward a target that does not respond to its own gradient steps. Every 500 steps we record the Bellman error on $\mathcal{D}$ (current network on both sides) and the value error $\frac{1}{|\mathcal{F}|}\sum_{(s,a)\in\mathcal{F}}|Q_{\theta}(s,a)-Q^{*}(s,a)|$ over all 4,795 feasible pairs. As a supervised baseline, OLS regression of $\log c$ on $\log k$ and $\log z$ is estimated on expanding windows.130130130Noise $\sigma=0.30$; windows from $n=10$ to 2,0002{,}000; held-out test set of 500 transitions.
@@ 7 Reinforcement Learning for Optimal Control
A handful of organizations have deployed reinforcement learning beyond simulation, achieving measurable gains on specific large-scale problems. Each deployment required substantial domain engineering and scientific tuning; these remain exceptional cases rather than standard practice. I review the most prominent field deployments, including ride-hailing dispatch at DiDi, data center cooling at Google, hotel revenue management, and financial order execution, before concluding with a simulation study on the bus engine replacement problem. In each case, the RL agent’s parameters were updated during a training phase conducted in simulation or on historical data, and the resulting policy was deployed with fixed weights (Section 2), with the exception of the hotel revenue management system, which learned in-field.
@@ 7.1 Ride-Hailing Dispatch
Qin et al. (2021) formalized DiDi’s order dispatching as a semi-Markov decision process131131131A semi-Markov decision process generalizes the standard MDP by allowing variable time between decisions. The discount factor $\gamma^{\tau_{k}}$ depends on the actual time elapsed $\tau_{k}$ rather than applying a fixed per-step discount. where each driver is an independent agent. The state of a driver consists of location (discretized into hexagonal zones) and time (bucketed into intervals). The action is the order assigned to the driver, with the option to remain idle. The reward is the trip fare. State transitions are determined by trip destinations, as completing an order transports the driver from origin to destination, changing their spatial state. The stochasticity arises from future demand, which determines the available actions at each state. The per-driver value function $V^{\pi}(s)$ represents the expected cumulative fare a driver can earn from state $s$ under dispatching policy $\pi$:
@@
where $\hat{r}_{i}$ is the predicted fare for order $i$, $\hat{\tau}_{i}$ is the estimated trip duration, and $s_{j}^{\prime}$ is the destination state. The value function is learned via temporal-difference methods from historical trip data. DiDi’s Cerebellar Value Network (CVNet; Tang et al., 2019) uses hierarchical coarse-coding132132132Coarse-coding represents a value function as a weighted sum over overlapping rectangular or hexagonal tiles that partition the state space (Sutton and Barto, 2018); each state activates the tiles that contain it, and a hierarchical variant stacks tiles at multiple resolutions to allow both coarse and fine generalization simultaneously. with a multi-resolution hexagonal grid, enabling transfer learning133133133Transfer learning initializes a model for a new task (a new city) with parameters trained on related tasks (existing cities), on the assumption that learned representations of traffic and demand patterns generalize across contexts and require less data to reach good performance in the new setting. across cities and robustness to data sparsity in low-traffic zones.
@@
Han et al. (2022) reported complementary results from Lyft, where the dispatching system optimizes driver assignment and repositioning jointly. Their value decomposition architecture134134134Value decomposition decomposes the platform’s global matching objective into individual driver value functions, each estimable via temporal-difference learning. The global optimum is recovered by optimizing the sum of individual values. assigns credit to individual driver decisions within the global objective. The production system demonstrated improvements in rider wait times and driver utilization, providing a second data point suggesting that similar approaches may transfer across platforms.
@@ 7.2 Hotel Revenue Management
Their system uses a two-step design. The RL agent observes the state $(t,s)$, where $t$ indexes the booking period within a $T=10$-period episode and $s$ is the average revenue per room sold to date, and selects an average discount level $a\in\{10\%,20\%,30\%\}$. A linear program then converts this scalar recommendation into a feasible capacity allocation across rate segments, accounting for each hotel’s channel preferences.135135135The RL component uses a modified on-policy Monte Carlo method with $\varepsilon$-greedy exploration, updating Q-values from realized returns after each completed episode, making it an in-field learner in the terminology of Section 2. This decomposition addresses practitioner acceptance, since managers understand a single discount recommendation, and circumvents the need to explicitly model demand arrivals or cancellation rates.
The field experiment randomly assigned five Hanting-brand hotels in Shanghai to the treatment group from ten candidates, with 271 additional Shanghai hotels serving as a donor pool for synthetic control estimation.136136136Synthetic control constructs a weighted combination of untreated hotels whose pre-treatment trend matches each treated hotel, avoiding the selection bias of simple before-after comparisons. Table 10 reports the average treatment effects over the pilot period (March–June 2015).
@@ 7.3 E-Commerce Dynamic Pricing
where $\text{uv}_{i,t}$ is unique visitors in period $t$ and $\tau$ is a reference lag.137137137DRCR normalizes revenue by traffic to remove demand fluctuations unrelated to pricing. The differencing further removes product-specific level effects, yielding a reward signal that is more concave than raw revenue and improves convergence stability.
Two algorithms are deployed: DQN for the discrete action formulation and DDPG for continuous actions. Both are pre-trained from demonstrations using historical specialist pricing actions (DQfD, DDPGfD) to address cold-start.138138138Pre-populating replay buffers with rollouts from heuristic or scripted policies is a general technique for reducing exploration burden in deep RL. In robotic grasping, Kalashnikov et al. (2018) collected 200,000 scripted grasping attempts at 15–30% success rate, sufficient to bootstrap a vision-based policy to 96% success; see Ibarz et al. (2021) for a survey of these and related warm-start strategies.139139139Standard A/B testing is infeasible because Chinese e-commerce regulations prohibit displaying different prices to different customers for the same product simultaneously. Liu et al. (2019) instead evaluate using difference-in-differences against “simi-products” (similar products not subject to algorithmic pricing) as controls.
@@ 7.4 Financial Order Execution
Nevmyvaka et al. (2006) applied tabular Q-learning to execution on real limit order book data from NASDAQ stocks.140140140The dataset covers 500 trading days of millisecond-level limit order book snapshots for AMZN, QCOM, and NVDA. The state space contains approximately 10,000 states; the horizon is $T=60$ seconds with discount $\gamma=1$. The state at time $t$ is $s_{t}=(t,q_{t},\psi_{t},\Delta_{t})$, where $t\in\{1,\ldots,T\}$ is time remaining, $q_{t}\in\{0,1,\ldots,Q\}$ is inventory remaining, $\psi_{t}$ is the discretized bid-ask spread, and $\Delta_{t}$ is the discretized signed volume imbalance.141141141Signed volume imbalance is the difference between buy and sell order volume near the best prices, normalized by total volume; positive imbalance typically predicts short-term price increases. Actions $a_{t}\in\{0,\delta,2\delta,\ldots,q_{t}\}$ specify shares to execute in the current interval. The reward is the negative per-period slippage contribution:
@@
where $p_{0}$ is the mid-quote at order arrival and $p_{t}^{\text{exec}}$ is the average execution price. Total implementation shortfall is $IS=-\sum_{t=1}^{T}r_{t}/Q$.142142142Implementation shortfall measures the total cost of executing a trade relative to the benchmark mid-quote at arrival. It captures market impact, timing cost, and opportunity cost of unexecuted shares. Because the objective is cost minimization, the Q-learning update uses $\min$ over next-period actions:
@@ 7.5 Supply Chain Inventory Management
Multi-echelon inventory systems coordinate ordering decisions across supply chain stages, where each stage’s order becomes the next stage’s incoming shipment. A retailer orders from a warehouse, which orders from a distribution center, which orders from a factory. The sequential nature creates complex dependencies, since an order placed upstream today affects downstream availability many periods later. The state space grows exponentially in the number of echelons, with $K$ stages each having $M$ possible inventory levels yielding $M^{K}$ states. Classical inventory theory provides closed-form solutions for special cases, most notably the echelon base-stock policy of Clark and Scarf (1960).
@@
Gijsbrechts et al. (2022) conducted a systematic evaluation of deep RL against classical base-stock policies across lost sales, dual sourcing, and multi-echelon configurations. The classical benchmark is the echelon base-stock policy, in which each stage $k$ maintains an echelon inventory position and orders to bring this position to a target level $S_{k}$.143143143Formally, the echelon inventory position at stage $k$ equals on-hand inventory at $k$ plus all inventory at downstream stages $1,\ldots,k-1$ plus in-transit inventory, minus backorders. The echelon base-stock policy of Clark and Scarf (1960) is optimal for serial systems with linear costs and backorders. For single-echelon systems with backorders, the optimal base-stock level is given by the newsvendor critical fractile $S^{*}=F_{D}^{-1}(b/(b+h))$. Table 13 summarizes results from their multi-echelon experiments.